# Deliberative XAI: How Explanations Impact Understanding and Decision-Making of AI Novices in Collective and Individual Settings

### v1

@@ 4. Results
Case 1: Group discusses changes of perspectives on the system’s deployment. Group B was composed of staff members and volunteers of a civil society organization. Three participants in this group changed their votes from ”Yes” in the first report to ”No” in the second report. In the excerpt below, B2 gives their perspective on the system’s deployment, and B3 responds with theirs, referring to explanations about the system’s features and weightings (System details, Base, B + Level 2, C), which fuel the discussion that follows:
@@
B2: I’m skeptical, but I’m still in favor of introducing it. Because it could be an aid and a relief for the staff working there.
@@
B3: I was originally in favor for these reasons, but since I’ve seen these parameters, I would be very much against it. Because I think there’s a lot of ideology in it. I think it’s no longer acceptable that men are favored over women and that duty of care only applies to women – this comes from a time that should be long gone.
B2: Those are strong arguments.
@@
B3: The things that come out are so absurd as well. For example, the apprenticeship of our Harald [the example job-seeker] was rated positively. He can’t use the apprenticeship at all for retraining. These things are given far too little consideration. […] As much as I like the idea, I don’t like the parameters.
B1: Did you vote yes first?
B3: I ticked yes at first, but I was really shocked at what was in there. […]
B1: What I’m wondering is, what would be the real benefit of introducing the system? What could be found out with much less effort? After all, employees should be able to see which degree you have quite quickly.
@@
B4: It’s a grid, a structure for the people who work at the agency so that they can quickly find a box.
The excerpt highlights how the provided information changed B3’s attitude toward the system and how this change, in turn, affected the group discussion. The combination of explanations and group discussion here had a major effect on participants’ perceptions, as three participants in this focus group changed their vote from ”yes” in the first self-report to ”no” in the second. Interactions like B2’s appreciation of B3’s arguments and B1’s questions about the system’s benefits depict processes of constructive group deliberation directly supported by the explanations.
Case 2: Group debates the meaning of ”unalterable” facts. Group D was composed of participants who were neither job-seekers nor domain experts and thus would be affected only indirectly by the system’s deployment. Even so, when D1 referred to an explanation describing how two different job-seekers would be judged by the system (System details, Base, C), it led to a discussion about the objectivity and adequacy of using data to represent a person:
@@
D3: I believe that the system can form the initial basis, based on the unalterable facts, which are of course weighted, but then it has to be enriched by a human being. […]
@@
D1: But I don’t believe that there are unalterable facts – well, not in this area. It’s all a question of representation and the lens through which you see the world.
@@
D3: When the job-seeker says, I only have four years of elementary school, then that’s four years of elementary school… […]
D2: That doesn’t mean that he can’t still be a very educated person.
D3: But that is hard to sell to an employer, right? […]
@@
D2: I’m skeptical about the data. You [D3] said the ”basis”, I think there are cracks in this basis. And I’m afraid of these cracks, that something will be pre-determined…
D3: But the human decision is always subjective.
@@
D2: That has to be weighed up. On the one hand, you have the arbitrariness of the individual employee, yes, and on the other hand you have an incomplete picture of a person.
D1: Or a false image.
D2: An incomplete one, I would say.
In this excerpt, the discussion surfaces discrepancies in how participants perceived the factuality and value of data. While the discussion in this focus group did not lead to a consensual decision in the time given, we still see this exchange as an important process in the deliberation on public AI systems. The merit of this discussion was later acknowledged by D1, who found the explanations confusing but stated that these exchanges were the ”centerpiece” and the most intriguing part. This highlights the role of explanations in the surfacing and debating of divergent opinions.
Case 3: Group discusses consequences of the system’s deployment for affected people. Group H was composed of staff members of a civil society organization specializing in helping young adults to take up apprenticeships. While all participants were clearly against the system’s deployment, the group decision task still led to a discussion about why specifically they opposed the system. The excerpt begins with H2’s comment about an explanation describing that job-seekers do not have a right of contestation (Context, Level 3, B):
H2: Just the fact that you can’t contest the decision, that’s a no-go.
H1: I think that having such systems introduced by institutions that themselves have very little idea of the real work – they may even want to help and support work systems, but that completely misses the mark because they are making decisions that are again driven by economic interests.
H4: And the life of the people who are then affected by the system is not represented at all, the psychosocial situation of the people is not taken into account.
H3: It is if they are depressed…
H4: Then it’s minus, minus, exactly. It simply deprives people of the opportunities they deserve. And that’s probably also my personal experience, I’ve worked with long-term unemployed adults who also had physical or mental impairments. The problem is that the longer I’m out of the job market, the fewer opportunities I have and people are already in a bad enough situation. And then you’re totally stuck in this [algorithmic] system. I just think that’s unfair.
H2: In terms of labor market policy, it may be justified, but in terms of social policy, it’s totally wrong.
The excerpt shows that even participants with unanimous attitudes toward the system found value in discussing and sharing their reasons for disapproving its deployment. H4 in particular appeared to be incensed by the system’s representation of soft factors. Their account of working with long-term job-seekers was later appreciated by the other participants, who stated that this perspective was valuable to them. The explanations in this case served to facilitate group conversation about the meaning and consequences of the system’s deployment and encouraged participants to share their views, stances, and experiences.
Summary. While Section 4.2.1 describes that explanations were effective in supporting group collaboration, this Section shows that their role in providing information for discussions between participants is equally important. Explanations helped participants in debating topics that are pivotal to the acceptance, trust, and application of ADM systems, in effect supporting their collaborative reasoning (Moshman and Geil, 1998). In this way, the explanations support what Kawakami et al. (2024) describe as ”deliberations around whether and under what conditions to move forward with developing or deploying” a public sector AI tool. This deliberation is important, as it allows participants to express and reflect on their opinions, feel heard, and consider other perspectives when making decisions.
@@ 5. Discussion
In this section, we discuss the implications of our findings for the design of explanations for AI novices (Section 5.1) as well as suggestions for their applications in collaborative settings (Section 5.2), which are summarized in Figure 7.
5.1. Reflections on the explanation design
Explanations improved or calibrated participants’ understanding, but measurements should be reconsidered. In Section 4.1.1, we described that while participants gave discrepant responses in their understanding self-reports (Table 3), they rarely felt that their understanding truly decreased or remained the same. Instead, we argue that the explanations changed their frame of reference for ”understanding” by revealing blind spots and making participants realize that their initial understanding was worse than assumed. Drawing from definitions in the cognitive sciences, the explanations thus supported participants in developing understanding by i) making participants realize that their initial understanding was incomplete (Rozenblit and Keil, 2002) and ii) supporting them in locating and closing relevant understanding gaps (Keil, 2006). The effect of participants reporting decreased perceived understanding following the explanations can be attributed to the fact that these explanations led them to attain a ”partial understanding” (Keil, 2019), but also showed them the difference to a complete, comprehensive understanding. The difficulty of self-assessing one’s understanding is a known issue (Rozenblit and Keil, 2002) and previous studies in XAI have shown similar effects caused by white-box explanations (Cheng et al., 2019), which, due to their high information density, led to increases in objective understanding but decreases in self-perceived understanding. This is not necessarily an issue regarding participants’ understanding, and together with the substantial increases in decision confidence, it can be seen as evidence that our explanations effectively improved and calibrated participants’ understanding. However, it also demonstrates that the relativity of internal cognitive states should be accounted for when measuring understanding. For this reason, recent work in XAI has proposed measurement based on abilities (Speith et al., 2024). This approach appears promising, considering that our single interview participants often found applying their knowledge in the study tasks helpful. Still, it likely requires modification when used in collective settings, as, for example, participants in focus groups never calculated the precise employment chance in task 2. The operationalization of the abilities-based approach for both individual and group settings is thus a direction for future research.
Participants considered all four explanation categories relevant and appreciated technical and contextual information. Our explanation design covered the categories data, system details, usage, and context. In the second self-reports, participants indicated which of these categories were ”helpful for their understanding” and ”influential to their decisions” (Section 3.3). In Section 4.1.5, we describe that focus groups reported that all categories helped their understanding but that system details more often than other categories influenced their decisions. In contrast, participants in single interviews reported all categories equally often as helpful and influential except data. Possible explanations for this effect include the forced prioritization due to the time limit, which led participants in single interviews to skip information about data as it was believed to be less relevant. In comparison, some focus groups proceeded chronologically through the explanations and thus spent more time on data as the first category. Further, the findings highlight that participants welcomed non-technical information about the system’s intention, ethical values, and practical usage (Figure 6) but also assigned high importance to the model’s exact feature weights. We see this as confirmation that participants’ information needs cover the entirety of an ADM system and that explanations for AI novices should aim to provide both technical (data, system details) and contextual (usage, context) information.
Explanations for decision subjects require a co-design approach. In Section 4.1.2, we describe that our explanations were most effective for participants who did not face challenges with either the language, large amounts of text, or the explanations’ structure. While these attributes are not entirely congruent with participants’ education levels, we observe that university-educated participants faced comparatively little difficulty interacting with the explanations, as the conversation excerpts in Sections 4.2.1+4.2.2 show. The differences were most apparent in Groups E and F, where participants stated that the explanations did not help them fully grasp the system as they required a high language and reading proficiency. We see this as evidence that a co-design approach similar to Weitz et al. (2024)’s work is required to build explanations for groups composed like Groups E and F, as it might be difficult to identify which content and format of explanations are required by decision subjects. On the other hand, the findings demonstrate the explanation design’s suitability for groups of domain experts, who also stated that this format of conveying information would be useful as a training measure.
Suggestions for improving the interaction design and combining individual and collective settings. In Sections 4.1.1 and 4.1.2, we described that according to participants’ feedback, the explanation design offers comprehensive and flexible information selection and encourages self-directed and active exploration, but that it also has a high access threshold and requires adjustment to the subdivided structure. For focus group participants, this design resulted in participants splitting and swapping explanations, sharing relevant parts, and discussing surprising information, which overall resembled an associative approach that supported group interaction but also led to a more fragmented picture of the system. In single interviews, participants could select and interact with information at their own pace and in a focused manner but, on the other hand, missed exchange and external perspectives. These effects can be described in the context of collaborative success and failure (Nokes-Malach et al., 2015), which states that groups require extra cognitive effort to coordinate and manage interactions, which can lead to memory limitations, information retrieval disruption, and blocking of thought processes if these coordination costs are too high. In contrast, individuals do not incur these cognitive costs, but of course can also not benefit from collaborative success mechanisms such as increased working memory or complementary knowledge. A future iteration of the explanation design could thus follow an individual with a collective interaction phase to combine the advantages and offset limitations. Further, a digital version would facilitate overview (summaries) and navigation (menus) while allowing for simple language options and cross-references. Lastly, the essential information of each explanation category, which in our design was contained in the base levels (Section 3.2), could use a balance of textual and visual design to ensure that essential information does not rely only on textual understanding.
On a side note, we emphasize that identifying explanation subtopics and splitting them up into levels of detail is challenging. Our explanations’ structure provides different levels of soundness and completeness (Kulesza et al., 2013), but deciding how the available information is allocated into this structure requires a subjective choice. In our study, detail levels 2 and 3 were supposed to convey more background rather than essential information gradually. Still, some participants stated that the most critical information for them resided in level 3. We see avenues for future work in selecting and hierarchically structuring information to be included in explanations for ADM systems, such that this structure allows for exploration while not obscuring essential or complex information.
5.2. Benefits and challenges of using explanations in collaborative settings
In Sections 4.2.1+4.2.2, we described that focus groups’ interactions with information led to ”collaborative success” (Nokes-Malach et al., 2015) when the group dynamic was constructive, i.e., argumentative (Mercier and Sperber, 2011), familiar (Johnson and Johnson, 1985), and solution-oriented (Niculae and Danescu-Niculescu-Mizil, 2016), and that explanations can inform conversations about central topics in public AI deployment (Kawakami et al., 2024). In the following, we discuss these findings by outlining the resulting benefits and challenges for collaborative XAI.
@@ 5.2. Benefits and challenges of using explanations in collaborative settings
Benefits: Explanations support collaborative reasoning, problem-solving, and deliberation. In the best cases, focus groups in our study had an open, communicative, and investigative atmosphere that facilitated the up-taking, sharing, and discussion of information. In these settings, the explanations showed their strengths by allowing for the distribution of reading tasks among group members, providing high levels of detail and breadth if needed, and offering different viewpoints that could be used as argumentative and conversational starting points. In this sense, the explanations fulfilled their aim of supporting collaboration and informed decision-making around deploying a public AI system (Kawakami et al., 2024) by creating common ground that participants used as a basis for deliberation. This deliberation process is the differentiating factor compared to ”one-to-one” (Naiseh et al., 2021) explanation settings, which might allow for more concentrated learning (Bloom, 1984), but miss the exchange of knowledge and perspectives that could become central for societal discourse about public AI (Züger and Asghari, 2023). While this collaborative atmosphere has several preconditions, such as containing cognitive biases (groupthink (Janis, 1972), equality bias (Naiseh et al., 2024)) and requiring trust between group members (Johnson and Johnson, 1985), our findings suggest that explanations can be an effective enabler in supporting these crucial group interactions. Regarding learning and understanding, XAI could further benefit from connecting to research showing that small groups can exhibit collective rationality (Moshman and Geil, 1998), outperform the wisdom of the crowds (Navajas et al., 2018), and solve tasks correctly even if no group member knows the solution beforehand (Smith et al., 2009).

### v2

@@ 4.1.5. Summary RQ1: Group and individual settings address different understanding facets in explainable AI
Group B was composed of staff members and volunteers of a civil society organization. Three participants in this group changed their votes from ‘Yes’ in the first report to ‘No’ in the second report. We found this change to be driven by three main deliberation elements (Stromer-Galley, 2007): ‘sourcing’ information, ‘reasoned arguments’ (opinion claims grounded in the information), and ‘engagement’ with the topic and between participants. In the excerpt, B2 and B3 weigh pros and cons of deployment. B3 grounds their arguments in explanations about the system’s features and weightings (system details), changing the discussion’s course:
@@
B2 I’m skeptical, but I’m still in favor of introducing it. Because it could be an aid and a relief for the staff working there.
@@
B3 I was originally in favor for these reasons, but since I’ve seen these parameters, I would be very much against it. Because I think there’s a lot of ideology in it. I think it’s no longer acceptable that men are favored over women and that duty of care only applies to women. This comes from a time that should be long gone.
B2 Those are strong arguments.
@@
B3 The things that come out are so absurd as well. For example, Harald’s apprenticeship was rated positively, but he can’t even use the apprenticeship for retraining. […] As much as I like the idea, I don’t like the parameters.
B1 Did you vote yes first?
B3 I ticked yes at first, but I was really shocked at what was in there [in the system]. […]
B1 What I’m wondering is, what would be the real benefit of introducing the system? […]
@@
B4 It’s a grid, a structure for the people who work at the agency, so that they can quickly find a box.
The excerpt highlights how the explanations led B3 to change their deployment decision and served as discussion triggers. In the resulting discussion, participants state both arguments (discrimination, what’s the benefit?) and opinions (disagreement with policy choices, AI can assist in decisions). Note that there is a difference between arguments (expression of reasoning processes that can be defended against critique) and opinions (expression of the speaker’s belief) (Stromer-Galley, 2007; Mercier and Landemore, 2012). While conceiving arguments to persuade interlocutors can result in confirmation bias (interpreting evidence such that it confirms existing beliefs) (Mercier and Sperber, 2011), the fact that B3 changed their attitude, in fact, indicates that the explanations acted against this bias. We argue that the excerpt thus shows a positive synergy in that the explanations provided grounds for ‘arguments’, which entered the discussion and provoked ‘collective reasoning’ and three decision swings. However, considering the large argumentative influence of B3, it should also be considered how the discussion would unfold if B3 had advocated for deployment. A case with comparable dynamics is described in 4.2.4.
4.2.3. Case 2 - disagreement: Group D debates normative positions regarding the algorithmic representation of people.
Group D was composed of participants who had been job-seeking in the past. When the group discussed the AI system’s deployment, the conversation shifted to how features that represent job-seekers’ profiles should be selected and weighed. This produced disagreement, an “important marker for deliberation” (Stromer-Galley, 2007) that displays heterogeneity of viewpoints, acts against polarization, and involves close examination of others’ reasoning. In the excerpt, D3 argues for the system’s deployment, while D1 argues against, and D2 acts as a mediator:
@@
D3 I believe that the system can form the initial basis, based on the unalterable facts, which are of course weighted, but then it has to be enriched by a human being. […]
@@
D1 But I don’t believe that there are unalterable facts – well, not in this area. It’s all a question of representation and the lens through which you see the world.
@@
D3 When the job-seeker says, ‘I only have four years of elementary school’, then that’s four years of elementary school… […]
D2 That doesn’t mean that he can’t still be a very educated person.
D3 But that is hard to sell to an employer, right? […]
@@
D2 I’m skeptical about the data. You [D3] said it’s the ‘basis’, I think there are cracks in this basis. And I’m afraid […] that something will be pre-determined…
D3 But the human decision is always subjective.
@@
D2 That has to be weighed up. On the one hand, you have the arbitrariness of the individual employee, yes, and on the other hand, you have an incomplete picture of a person.
D1 Or a false image.
D2 An incomplete one, I would say.
The group here discusses diverging views and expresses opinions. While these opinions are meant to persuade and defend, they are grounded in lived experiences rather than in the explanations. The central conflict develops between D1’s belief that the system misrepresents reality and D3’s viewpoint that it can increase objectivity and assist in decisions. The discussion here did not lead to a consensus on the deployment decision in the given time, but resulted in a majority vote for ‘No’. We argue that it still illustrates an important process in the deliberation on public AI systems: Participants again ‘sourced’ information that was turned into arguments, but the debate led to a more fundamental topic that surfaced discrepancies which would impede finding a collective decision. The fact that participants then engaged in ‘disagreement’ is a sign of productive deliberation, as it shows that there were diverse viewpoints, that no polarization or ‘groupthink’ (Janis, 1971) occurred, and that the proposal was closely examined based on the information given (Stromer-Galley, 2007). In a real setting, this form of debate could serve as a fruitful basis to investigate whether the system is in the ‘public interest’ (Züger and Asghari, 2023) and to host ‘early-stage deliberations’ (Kawakami et al., 2024) on the system during development. The merit of this debate was further later acknowledged by D1, who found the explanations confusing but stated that these exchanges were the study’s “centerpiece” and most intriguing part. We argue that the interplay between explanations and group discussion here supported a (simulated) evidence-informed policy-making process (Mair et al., 2019).
4.2.4. Case 3 - groupthink? Group A follows a minority position and votes for system deployment
Group A was composed of volunteers from a civil society organization. Three participants in this group changed their decisions, shifting from ‘No’ to ‘Yes’ after the explanation phase. We explain these changes with three aspects: First, Group A focused on the explanation category data and did not interact much with other categories (Figure 6). This meant less attention was paid, for example, to the system’s feature selection and weightings that were decisive in Cases 1 and 2. Second, participants of Group A stated that they were not directly affected by the system, as they were retired, implying low ‘engagement’: “It doesn’t affect me anymore and I think to myself, yeah…” (A1). Third, participants prioritized group concurrence above a “careful, critical scrutiny” (Janis, 1971). The following excerpt illustrates the tipping point for the collective decision:
A3 I still think the system is better, even if there are still mistakes in it, than sitting opposite someone [a counselor] who doesn’t like you… […]
A2 So rather ‘no’?
A1 Yes, as A3 says, it’s… I don’t know.
A3 Yes and no… […]
A2 I mean, it can’t be avoided, it will happen. I’m convinced of that, whether we like it or not, it’s done.
A1 It won’t affect us anymore, at least not in the employment office. […] I agree with the majority.
A2 But that’s difficult now.
A3 I’ll stick with ‘yes’. My daughter would say I shouldn’t think so negatively, especially with AI. […]
A2 I say ‘yes’ too. […] You, A1 and A4, can tip the scales.
A1 I say ‘yes’ now too, but not because I’ve changed my mind, but because I want an overall solution.
A4 I say ‘yes’ but I’m leaning towards ‘no’.
Despite articulated reservations, all participants ultimately decided to vote for deployment. We compared this excerpt with characteristics of ‘groupthink’, a “mode of thinking” in which people value concurrence higher than consideration of alternative courses of action (Janis, 1971). This mode produces defective decision-making processes due to three key aspects: strong social identification with the group, salient norms, and a perceived low self-efficacy to make the decision (Baron, 2005). The excerpt clearly demonstrates two of these aspects: A1 changes their decision due to a desire for group harmony, and A4 follows suit (group identification). Further, both A1 and A3 express their uncertainty and sway between options (low self-efficacy). While A2’s statement that AI is inevitable is an opinion rather than an argument (neither ‘sourced’ nor the product of evident reasoning), it triggers the group to make a quick decision that disregards any remaining ‘disagreement’. Although the process can be connected to aspects of ‘groupthink’ (Janis, 1971), such as rationalizations of flawed logic and self-censorship, these aspects are not nearly as pronounced as in the literature (Janis, 1971, 1972; Baron, 2005). For example, the group did not share an illusion of unanimity, and the uncertainty among participants suggests no guiding salient norms. Still, as participants avoided ‘disagreement’ and instead followed decisions of other, the excerpt presents a suboptimal deliberation process (Baron, 2005). In part, this can be attributed to the explanation’s failure to make all fundamental information easily available and to not encourage analytical thinking over intuitive, heuristical thinking (Buçinca et al., 2021). In addition, the group might have missed a role that explicitly takes the opposing viewpoint to fuel discussion, which was identified to benefit deliberation in previous XAI research (Chiang et al., 2024). The implications of these findings are discussed in Section 5.
4.2.5. Summary RQ2
The findings in this Section demonstrate how explanations supported deliberation in focus groups and single interviews. Many participants reported improved decision confidence and changed their deployment decisions based on the explanations, often due to a disillusionment regarding the AMS algorithm’s assumed merits. These changes occurred in both settings, suggesting that the explanations supported public and internal deliberation (Mercier and Landemore, 2012). In group settings, participants used explanations when discussing deployment, as illustrated in Case 1. Case 2 further highlights that explanations surfaced discrepancies in personal beliefs and produced productive conflict. In contrast, Case 3 shows a deployment decision based more on concurrence-seeking than on ‘collaborative reasoning’ (Moshman and Geil, 1998). However, we hesitate to label the exchange as ‘groupthink’, as it does not align with all factors that characterize the phenomenon (Baron, 2005). Based on these findings, we argue that explanations can support people in considering if AI systems are in the public interest and to discuss “whether and under what conditions to move forward with developing or deploying” them (Züger and Asghari, 2023). To achieve this, both the explanations and the group setting need to i) be designed so that they allow for the easy sourcing of information for arguments, ii) make all relevant information available as soon as possible, and iii) include mechanisms that encourage participants to examine both the proposal and their positions closely. Matching explanations and social setting to support ‘elements of deliberation’ (Stromer-Galley, 2007) thus presents promising starting points for future research on how explainable AI can promote public deliberation on AI.
@@ 5. Discussion
In this section, we discuss how our findings answer our two main research questions: Whether a question-driven, modular explanation design supports AI novices’ understanding in groups and individual settings (RQ1) and how AI novices used these explanations to deliberate about AI systems (RQ2). We describe the advantages of both social settings for explainable AI, outline which real-world use cases would benefit from our explanation design, discuss whether the explanations improved understanding, and provide suggestions for their design improvements. We summarize the implications of our findings in Figure 8.
5.1. Do AI novices learn and deliberate about AI better together or individually?
@@
In Section 4.1, we described that explanations produced shared understanding in groups, involving both cognitive and social mechanisms of “collaborative success” (Nokes-Malach et al., 2015). Section 4.2 further showed that explanations improved participants’ decision confidence and provided grounds for different elements of deliberation (Stromer-Galley, 2007), such as reasoned arguments and disagreement. In the best cases, focus groups in our study had a familiar (Johnson and Johnson, 1985) and solution-oriented (Niculae and Danescu-Niculescu-Mizil, 2016) atmosphere that facilitated sharing and discussing information. In these settings, the modular explanation structure showed its strengths by allowing for the distribution of tasks among group members, providing high levels of detail and breadth if needed, and offering different viewpoints that could be used as argumentative and conversational starting points. In this sense, the explanations fulfilled their aim of supporting learning and deliberation about a public AI system (Kawakami et al., 2024). The interaction between group members is the differentiating factor compared to “one-to-one” (Naiseh et al., 2021) explanation settings. In our study, single interviews allowed for more focused engagement with explanations and a form of “internal deliberation” (Mercier and Landemore, 2012) but lacked the exchange of knowledge and perspectives with others that is deemed central for deliberation about public AI (Züger and Asghari, 2023). Regarding learning and deliberation, XAI would thus benefit from researching how group settings can be used to leverage collective reasoning (Moshman and Geil, 1998), wisdom of the crowds (Navajas et al., 2018), and performance increases through peer discussion (Smith et al., 2009). However, the benefits of group settings have several preconditions, such as the containment of cognitive biases (groupthink (Janis, 1971), equality bias (Naiseh et al., 2024)) and, crucially, a trusting social dynamic (Chiang et al., 2023).